(NOTE: I got a call from Shiva today and our LANRover that was sent in for repairs was found to be in good-working order; that is, they could find nothing wrong with it. --Bob)
- Skyline update. Justin, Carm and Garnet and exploring the use of this network monitoring software and are making progress.
- Library CD-ROM server. Dave nor Bob have heard anything about this. Bob did hear from Geri Linkins that she has not seen a purchase requisition for the memory that needs to be ordered.
- Key systems (telephones). Manola is working on this and calling offices to check on where these sets are located.
- Bowman setup. Carm, Garnet and Joan have set up Bowman on ethernet and all of the Mac SEs have been moved into place there.
- 2400 baud modems. Still getting reports about the 2400 baud modems not answering. Carm will chek them again and remove any that are faulty. Bob will check with Jack to see if we can take his Zyxel modem and add it to our 2400 baud modem pool (NOTE: Jack is willing to do this --Bob)
- TV "banner" for the InfoChannel. Carm found out that you must use 2 synchronous signals to do this. He thinks we might be able to use our Chyron titler to do this if it still works. He will check on this.
- SUNYNet will drop its Internet gateway on Sunday 2/19 from 8am - Noon.
- Bob asked how we might quickly program switchboard to handle calls about cancellations during inclement weather. Brenda pointed out that we handle cancellations as one of the menu items of the Campus Activities Hotline (x4444). Bob will talk to Paula about updating x4444 if there are campus/class cancellations. (NOTE: Paula is working on this now. --Bob) - Cabletron meeting. Bob, Carm and Garnet met with our Cabltron sales rep, Sharon O'Neill. She wants to send an engineer here to look at our residence halls and make a proposal for wiring them for networking. She understands that even though she will do this at no cost to us, we may not go with Cabletron. She's willing to take that risk (Way to go, Sharon!)
- Fix the LANRovers! (Bob/Justin)
- Fix the 2400 baud modems. Install Jack's Zyxel modem in the pool. (Carm)
- Update AUDIX/Switch records. (Brenda)
- check on progress being made to install the backup generator for the switch. (Bob)
- Bob presented his revised "SUNYNet Alert Action Plan". There are still 3 questions that need to be asked--but they are the same 3 questions that we will need to ask for each problem we want to alert the campus about:
- Bob asked everyone again to submit their plans via email so we could begin to review them prior to our next (2/21) team meeting.
(Dave/Justin--you need to get together on this ASAP!)
5. Next meeting (Tues, 2/21 @ 8:30am) --> **Please** arrive a few minutes early so we can start promptly @ 8:30am.